OpenStudy (anonymous):

The point (−3, 2 ) lies on a circle whose equation is ( x + 3) + ( y + 1) = r 2 . Which of the following must be the radius of the circle? 3 A B 10 C 9 10 D

OpenStudy (anonymous):

You mean (x + 3)^2 + (y+1)^2 = r^2, right? Plug in the values for x and y, to get: 0^2 + 3^2 = r^2 -> r^2 = 9 -> r = 3, because the radius must be positive.

OpenStudy (anonymous):

a

OpenStudy (anonymous):

no... thats the equation that was on my math pg

OpenStudy (anonymous):

thnx Mit :)